Solve the equation for y. 7x - 3y = 4
Answer by Nihal@SriLanka(22)

You cannot actually solve this equation for y - all you can do is express y using the other terms
Subtracting 7x from both sides of the equation we have

7x-3y-7x = 4-7x
i.e. -3y = 4-7x
i.e. -y = (4-7x)/3
i.e. y = (7x-4)/3 which is how you could express y using the other terms.

If you need further clarification on what is happening here you may email sumanapala@gmail.com.